### v1

@@ 3 The Na and O distribution in GES OCs
1 | NGC 6253 | 3.24 | 68 | 0.478 | 0.099 | 0.38 | 0.36 | 0.13 |
2 | M 67 | 4.27 | 38 | 0.040 | 0.054 | 0.05 | 0.02 | 0.10 |
3 | NGC 3532 | 0.40 | 36 | 0.000 | 0.078 | 0.02 | 0.01 | 0.12 |
4 | Trumpler 5 | 4.27 | 26 | -0.245 | 0.038 | 0.23 | 0.35 | 0.06 |
5 | NGC 2141 | 1.86 | 23 | 0.004 | 0.064 | 0.01 | 0.05 | 0.09 |
6 | Ruprecht 134 | 1.66 | 20 | 0.399 | 0.045 | 0.41 | 0.27 | 0.09 |
7 | Trumpler 20 | 1.86 | 26 | 0.184 | 0.041 | 0.17 | 0.13 | 0.08 |
8 | NGC 2243 | 4.37 | 18 | -0.279 | 0.041 | 0.28 | 0.47 | 0.06 |
9 | Berkeley 32 | 4.90 | 17 | -0.141 | 0.085 | 0.16 | 0.29 | 0.12 |
10 | NGC 2516 | 0.24 | 15 | -0.039 | 0.051 | 0.05 | 0.05 | 0.08 |
11 | NGC 2158 | 1.55 | 14 | -0.076 | 0.038 | 0.05 | 0.16 | 0.09 |
12 | NGC 2425 | 2.40 | 14 | -0.059 | 0.066 | 0.07 | 0.14 | 0.12 |
13 | NGC 6067 | 0.13 | 13 | 0.216 | 0.141 | 0.23 | 0.03 | 0.15 |
14 | Berkeley 21 | 2.14 | 13 | -0.204 | 0.068 | 0.20 | 0.21 | 0.06 |
15 | NGC 6005 | 1.26 | 13 | 0.343 | 0.018 | 0.31 | 0.22 | 0.08 |
16 | Berkeley 81 | 1.15 | 13 | 0.348 | 0.057 | 0.36 | 0.25 | 0.14 |
17 | NGC 6259 | 0.34 | 12 | 0.379 | 0.045 | 0.38 | 0.17 | 0.10 |
18 | Haffner 10 | 3.80 | 12 | -0.079 | 0.047 | 0.07 | 0.12 | 0.10 |
19 | NGC 2477 | 1.12 | 11 | 0.156 | 0.072 | 0.21 | 0.13 | 0.15 |
20 | Trumpler 23 | 0.71 | 11 | 0.346 | 0.079 | 0.37 | 0.21 | 0.04 |
21 | Blanco 1 | 0.10 | 11 | -0.083 | 0.065 | 0.07 | 0.05 | 0.09 |
22 | NGC 2420 | 1.74 | 37 | -0.098 | 0.088 | 0.08 | 0.16 | 0.11 |
23 | NGC 6705 | 0.31 | 10 | 0.282 | 0.023 | 0.28 | 0.06 | 0.04 |
24 | NGC 6802 | 0.66 | 10 | 0.269 | 0.029 | 0.27 | 0.15 | 0.06 |
25 | 47 Tuc (GC) | 66 | 0.44 | 0.20 | 0.43 | 0.75 | 0.30 |
@@ A.4 Comparison of GES and RVs
Notes: 1 Ebrahimi et al. (2022); 2 Almeida et al. (2023) |

### v2

@@ 3 The Na and O distribution in GES OCs
| 1 | NGC 6253 | 3.24 | 68 | 0.478 | 0.099 | 0.38 | 0.36 | 0.13 |
| 2 | M 67 | 4.27 | 38 | 0.040 | 0.054 | 0.05 | 0.02 | 0.10 |
| 3 | NGC 3532 | 0.40 | 36 | 0.000 | 0.078 | 0.02 | 0.01 | 0.12 |
| 4 | Trumpler 5 | 4.27 | 26 | -0.245 | 0.038 | 0.23 | 0.35 | 0.06 |
| 5 | NGC 2141 | 1.86 | 23 | 0.004 | 0.064 | 0.01 | 0.05 | 0.09 |
| 6 | Ruprecht 134 | 1.66 | 20 | 0.399 | 0.045 | 0.41 | 0.27 | 0.09 |
| 7 | Trumpler 20 | 1.86 | 26 | 0.184 | 0.041 | 0.17 | 0.13 | 0.08 |
| 8 | NGC 2243 | 4.37 | 18 | -0.279 | 0.041 | 0.28 | 0.47 | 0.06 |
| 9 | Berkeley 32 | 4.90 | 17 | -0.141 | 0.085 | 0.16 | 0.29 | 0.12 |
| 10 | NGC 2516 | 0.24 | 15 | -0.039 | 0.051 | 0.05 | 0.05 | 0.08 |
| 11 | NGC 2158 | 1.55 | 14 | -0.076 | 0.038 | 0.05 | 0.16 | 0.09 |
| 12 | NGC 2425 | 2.40 | 14 | -0.059 | 0.066 | 0.07 | 0.14 | 0.12 |
| 13 | NGC 6067 | 0.13 | 13 | 0.216 | 0.141 | 0.23 | 0.03 | 0.15 |
| 14 | Berkeley 21 | 2.14 | 13 | -0.204 | 0.068 | 0.20 | 0.21 | 0.06 |
| 15 | NGC 6005 | 1.26 | 13 | 0.343 | 0.018 | 0.31 | 0.22 | 0.08 |
| 16 | Berkeley 81 | 1.15 | 13 | 0.348 | 0.057 | 0.36 | 0.25 | 0.14 |
| 17 | NGC 6259 | 0.34 | 12 | 0.379 | 0.045 | 0.38 | 0.17 | 0.10 |
| 18 | Haffner 10 | 3.80 | 12 | -0.079 | 0.047 | 0.07 | 0.12 | 0.10 |
| 19 | NGC 2477 | 1.12 | 11 | 0.156 | 0.072 | 0.21 | 0.13 | 0.15 |
| 20 | Trumpler 23 | 0.71 | 11 | 0.346 | 0.079 | 0.37 | 0.21 | 0.04 |
| 21 | Blanco 1 | 0.10 | 11 | -0.083 | 0.065 | 0.07 | 0.05 | 0.09 |
| 22 | NGC 2420 | 1.74 | 37 | -0.098 | 0.088 | 0.08 | 0.16 | 0.11 |
| 23 | NGC 6705 | 0.31 | 10 | 0.282 | 0.023 | 0.28 | 0.06 | 0.04 |
| 24 | NGC 6802 | 0.66 | 10 | 0.269 | 0.029 | 0.27 | 0.15 | 0.06 |
| 25 | 47 Tuc (GC) | 66 | 0.44 | 0.20 | 0.43 | 0.75 | 0.30 |
@@ A.4 Comparison of GES and RVs
| Notes: 1 Ebrahimi et al. (2022); 2 Almeida et al. (2023) | ||||||||||